What is the P-value for the testing of the following hypothesis for someone who gets 25 right out of 100? 30 right? 35 Right? 50 right?
The null hypothesis is p=.2 that a person is guessing in an ESP experiment where one person randomly draws one of 5 cards and the "reciever" person tries to determine which card it is. The alternative hypothesis is p>.2 is to show support that the person is doing better than guessing.

Answer by stanbon(48568)

What is the P-value for the testing of the following hypothesis for someone who gets 25 right out of 100? 30 right? 35 Right? 50 right?
The null hypothesis is p=.2 that a person is guessing in an ESP experiment where one person randomly draws one of 5 cards and the "reciever" person tries to determine which card it is. The alternative hypothesis is p>.2 is to show support that the person is doing better than guessing.
------------------------
Ho: p=0.2
Ha: p>0.2
--------------
If p-hat = 25/100 = 0.25
then z(0.25)= (0.25-0.20)/[sqrt[(0.2*0.8/25)] = 0.05/0.04/5 = 5*0.05/0.04
= 0.625
The the p-value associated with this z value is 0.266
--------------
Do the same for p-hat = 30/100, then p-hat=35/100, the p-hat=50/100
=========================
Cheers,
Stan H.